OpenStudy (anonymous):

When \frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25} is computed and written as a mixed number in simplest form, what is the fractional part of that mixed number?

OpenStudy (anonymous):

Can you use the equation tool? I can't read it.

OpenStudy (anonymous):

HOW DO YOU USE THE EQUATION BAR

OpenStudy (tkhunny):

you are SO CLOSE. Just include the code in \ ( code \ ) tags. \( \frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25}\)

Elsa213 (elsa213):

\(\frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25}\) *

OpenStudy (anonymous):

first rearrange the equation